The length of time a full length movie runs from opening to credits is normally distributed with a mean of 1.9 hours and standard deviation of 0.3 hours. Calculate the following:

  1. A random movie is between 1.5 and 2.0 hours.
  2. A movie is longer than 2.4 hours.
  3. The length of movie that is shorter than 94% of the movies.
1
Expert's answer
2021-06-23T17:46:23-0400

Solution:

Given, "\mu=1.9,\sigma=0.3"

"X\sim N(\mu,\sigma)"

(i):

"P(1.5<X<2)=P(X<2)-P(X<1.5)
\\=P(z<\dfrac{2-1.9}{0.3})-P(z<\dfrac{1.5-1.9}{0.3})
\\=P(z<0.33)-P(z<-1.33)
\\=P(z<0.33)-[1-P(z<1.33)]
\\=0.62930-[1-0.90824]
\\=0.53754"

(ii):

"P(X>2.4)=1-P(X\le2.4)
\\=1-P(z\le \dfrac{2.4-1.9}{0.3})
\\=1-P(z\le 1.67)
\\=1-0.95254
\\=0.04746"

(iii):

"P(X<x)=0.94
\\\Rightarrow P(z<\dfrac{x-1.9}{0.3})=0.94
\\\Rightarrow \dfrac{x-1.9}{0.3}=1.55
\\\Rightarrow x=2.365"

Required length of movie is 2.365 hours.